The state education commission wants to estimate the fraction of tenth grade students that have reading skills at or below the eighth grade level.

Step 2 of 2 : Suppose a sample of 323
tenth graders is drawn. Of the students sampled, 256
read above the eighth grade level. Using the data, construct the 80%
confidence interval for the population proportion of tenth graders reading at or below the eighth grade level. Round your answers to three decimal places.
Answer #1

We have given,                      
                      
x=323-256=67                  
n=323                  
Estimate for sample proportion=   0.2074          
Level of significance is =1-0.80=0.2              
Z critical value(using Z table)=1.282              
                      
Confidence interval formula is                      


=(0.179,0.236)
                  
Lower limit for confidence interval=0.179          
                      
Upper limit for confidence interval=0.236